OpenStudy (anonymous):

The function g(t) = 2t represents the number of guitar lessons, g(t), you can complete in t months. How many guitar lessons can you complete in 7 months?

Directrix (directrix):

Given that t = 7, then g(7) = 2 (7) = 14.

OpenStudy (hoblos):

g(t) = 2t for t=7 g(t) = 2*7 =14